OpenStudy (anonymous):

Simplify the expression 10(x-y)-4(x+y)

OpenStudy (anonymous):

use distributive property: a(x+y) = ax+ay ax +bx = (a+b)x

OpenStudy (anonymous):

distributive property. 10(x-y)-4(x+y) 10x-10y-4x-4y 6x-14y

OpenStudy (nikita2):

0, if x=0 and y = 0